# A Generalized Analytical Framework for the Nonlinear Best-Worst Method

### v1

@@ 4.3 Numerical Examples
Example 1: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$, $\epsilon_{4}=0.5363$, $\epsilon_{4,2}=0.5163$ and $\epsilon_{4,3}=0.5422$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.5422$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=8.4987$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4868$. Therefore, $w_{1}^{*}=[0.4855,0.4868]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0549,0.0574]$, $w_{3}^{*}=[0.1975,0.1981]$, $w_{4}^{*}=[0.2024,0.2029]$ and $w_{5}^{*}=[0.0571,0.0573]$.
@@
[…]
Table 3 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 2: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$, $\epsilon_{4}=1.7228$, $\epsilon_{4,2}=1.7882$ and $\epsilon_{4,3}=1.7999$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=1.7999$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=14.8760$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.6205$. Therefore, $w_{1}^{*}=[0.6200,0.6205]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0429,0.0437]$, $w_{3}^{*}=[0.1607,0.1609]$, $w_{4}^{*}=[0.1340,0.1341]$ and $w_{5}^{*}=[0.0416,0.0417]$.
@@
[…]
Table 4 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 3: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$, $\epsilon_{4}=0.8086$, $\epsilon_{4,2}=0.8825$ and $\epsilon_{4,3}=0.8975$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.8975$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=5.4354$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4280$. Therefore, $w_{1}^{*}=[0.4269,0.4280]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0842,0.0866]$, $w_{3}^{*}=[0.1831,0.1836]$, $w_{4}^{*}=[0.2250,0.2255]$ and $w_{5}^{*}=[0.0785,0.0787]$.
@@
[…]
Table 5 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 4: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$, $\epsilon_{4}=0.4865$, $\epsilon_{4,2}=0.0.5480$ and $\epsilon_{4,3}=0.5458$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,2}=0.5480$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,2})\times(a_{45}+\epsilon_{4,2})=6.8344$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4077$. Therefore, $w_{1}^{*}=[0.4074,0.4077]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.1558,0.1560]$, $w_{3}^{*}=[0.2407,0.2413]$, $w_{4}^{*}=[0.1359,0.1360]$ and $w_{5}^{*}=[0.0596,0.0597]$.
@@
[…]
Table 6 shows that the obtained interval-weights, $\epsilon^{*}$ and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 5: Here, $D_{1}=\{2\}$, $D_{2}=\{3\}$ and $D_{3}=\phi$. By equations (13) and (16), we get $\epsilon_{2}=0.3028$, $\epsilon_{3}=0.4384$ and $\epsilon_{2,3}=0.5$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{2,3}\}=\epsilon_{2,3}=0.5$. Now, equation 34 gives $\tilde{a}_{14}^{*}=(a_{12}+\epsilon_{2,3})\times(a_{24}+\epsilon_{2,3})=2.25$. From equation (20), we get
@@
[…]
@@
Example 6: Here, $n_{1}=2$, $n_{2}=1$, $D_{1}=\{3\}$ and $D_{2}=D_{3}=\phi$. By equations (13), we get $\epsilon_{2}=0.1623$. So, from equation (41), it follows that $\epsilon^{*}=\epsilon_{2}=0.1623$. Now, equation 42 gives $\tilde{a}_{14}^{*}=\tilde{a}_{24}^{*}=(a_{13}+\epsilon_{2})\times(a_{34}+\epsilon_{2})=6.8378$. From equation (43), we get
@@
[…]
Note that both $c_{1}$ and $c_{2}$ have the same optimal weights.
@@ 5 Conclusions and Future Directions
The BWM has emerged as a powerful MCDM tool, widely adopted in real-world applications. This study introduces a generalized analytical framework for its nonlinear model, compatible with any scale and any number of DMs. The framework derives closed-form solutions for optimal interval-weights, CI and CR, while also identifying the best weight set from all possible solutions. In the presence of multiple best/worst criteria, a key modification in the original optimization model ensures the selection of optimal weight sets in which criteria with equal preferences receive the same weights. The research advances BWM methodology by enhancing its capability to handle both unique and multiple optimal solutions across all preference values. By eliminating the need for specialized optimization software, the framework significantly improves computational efficiency. It also rectifies previously reported inaccuracies in CI calculations and transforms CR into an effective real-time consistency measure for immediate DM feedback.
Several important challenges remain unresolved in this research. Specifically, analytical solutions for optimal weights have not yet been developed for certain BWM variants, including the Euclidean BWM [22] and $\alpha$-FBWM [30].

### v2

@@ 4.3 Numerical Examples
Example 1: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equations (13) and (16), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$, $\epsilon_{4}=0.5363$, $\epsilon_{4,2}=0.5163$, and $\epsilon_{4,3}=0.5422$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.5422$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=8.4987$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4868$. Therefore, $w_{1}^{*}=[0.4855,0.4868]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0549,0.0574]$, $w_{3}^{*}=[0.1975,0.1981]$, $w_{4}^{*}=[0.2024,0.2029]$, and $w_{5}^{*}=[0.0571,0.0573]$.
@@
is the best optimally modified PCS. Thus, by equation (8), $\{0.4857,0.0571,0.1976,0.2024,0.0572\}$ is the best optimal weight set. Now, by equation (6), we get CR =0.1037=0.1037.
Table 3 shows that the obtained $\epsilon^{*}$, interval-weights, and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 2: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equations (13) and (16), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$, $\epsilon_{4}=1.7228$, $\epsilon_{4,2}=1.7882$, and $\epsilon_{4,3}=1.7999$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=1.7999$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=14.8760$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.6205$. Therefore, $w_{1}^{*}=[0.6200,0.6205]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0429,0.0437]$, $w_{3}^{*}=[0.1607,0.1609]$, $w_{4}^{*}=[0.1340,0.1341]$, and $w_{5}^{*}=[0.0416,0.0417]$.
@@
is the best optimally modified PCS. Thus, by equation (8), $\{0.6200,0.0436,0.1607,0.1340,0.0417\}$ is the best optimal weight set. Now, by equation (6), we get CR =0.1663=0.1663.
Table 4 shows that the obtained $\epsilon^{*}$, interval-weights, and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 3: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equations (13) and (16), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$, $\epsilon_{4}=0.8086$, $\epsilon_{4,2}=0.8825$, and $\epsilon_{4,3}=0.8975$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,3}=0.8975$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,3})\times(a_{45}+\epsilon_{4,3})=5.4354$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4280$. Therefore, $w_{1}^{*}=[0.4269,0.4280]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.0842,0.0866]$, $w_{3}^{*}=[0.1831,0.1836]$, $w_{4}^{*}=[0.2250,0.2255]$, and $w_{5}^{*}=[0.0785,0.0787]$.
@@
is the best optimally modified PCS. Thus, by equation (8), $\{0.4270,0.0861,0.1832,0.2251,0.0786\}$ is the best optimal weight set. Now, by equation (6), we get CR =0.3119=0.3119.
Table 5 shows that the obtained $\epsilon^{*}$, interval-weights, and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 4: Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equations (13) and (16), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$, $\epsilon_{4}=0.4865$, $\epsilon_{4,2}=0.0.5480$, and $\epsilon_{4,3}=0.5458$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{4},\epsilon_{4,2},\epsilon_{4,3}\}=\epsilon_{4,2}=0.5480$. Now, equation 34 gives $\tilde{a}_{15}^{*}=(a_{14}+\epsilon_{4,2})\times(a_{45}+\epsilon_{4,2})=6.8344$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.4077$. Therefore, $w_{1}^{*}=[0.4074,0.4077]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.1558,0.1560]$, $w_{3}^{*}=[0.2407,0.2413]$, $w_{4}^{*}=[0.1359,0.1360]$, and $w_{5}^{*}=[0.0596,0.0597]$.
@@
is the best optimally modified PCS. Thus, by equation (8), $\{0.4076,0.1559,0.2409,0.1360,0.0596\}$ is the best optimal weight set. Now, by equation (6), we get CR =0.1471=0.1471.
Table 6 shows that the obtained $\epsilon^{*}$, interval-weights, and the best optimal weight set coincide with their actual values, validating the proposed framework.
Example 5: Here, $D_{1}=\{2\}$, $D_{2}=\{3\}$, and $D_{3}=\emptyset$. By equations (13) and (16), we get $\epsilon_{2}=0.3028$, $\epsilon_{3}=0.4384$, and $\epsilon_{2,3}=0.5$. So, from equation (33), it follows that $\epsilon^{*}=\max\{\epsilon_{2},\epsilon_{3},\epsilon_{2,3}\}=\epsilon_{2,3}=0.5$. Now, equation 34 gives $\tilde{a}_{14}^{*}=(a_{12}+\epsilon_{2,3})\times(a_{24}+\epsilon_{2,3})=2.25$. From equation (20), we get
@@
Similarly, ${w_{1}^{u}}^{*}=0.36$. Therefore, $w_{1}^{*}=[0.36,0.36]$. Computing in the same fashion, we obtain $w_{2}^{*}=[0.24,0.24]$, $w_{3}^{*}=[0.24,0.24]$, and $w_{4}^{*}=[0.16,0.16]$. Therefore, $\{0.36,0.24,0.24,0.16\}$ is the only optimal weight set. Now, by equation (6), we get CR =1=1.
@@
Example 6: Here, $n_{1}=2$, $n_{2}=1$, $D_{1}=\{3\}$, and $D_{2}=D_{3}=\emptyset$. By equation (13), we get $\epsilon_{2}=0.1623$. So, from equation (41), it follows that $\epsilon^{*}=\epsilon_{2}=0.1623$. Now, equation 42 gives $\tilde{a}_{14}^{*}=\tilde{a}_{24}^{*}=(a_{13}+\epsilon_{2})\times(a_{34}+\epsilon_{2})=6.8378$. From equation (43), we get
@@
Similarly, ${w_{1}^{u}}^{*}={w_{2}^{u}}^{*}=0.3833$. Therefore, $w_{1}^{*}=w_{2}^{*}=[0.3833,0.3833]$. Computing in the same fashion, we obtain $w_{3}^{*}=[0.1773,0.1773]$, and $w_{4}^{*}=[0.0561,0.0561]$. Therefore, $\{0.3833,0.3833,0.1773,0.0561\}$ is the only optimal weight set. Now, by equation (6), we get CR =0.0436=0.0436.
Note that for $c_{1}$ and $c_{2}$, both the interval-weights and the best weights coincide. Additionally, Table 7 indicates that both optimization models (3) and (37) achieve the same optimal objective value. This suggests that optimization model (37) has selected the optimal weight sets in which the weights for $c_{1}$ and $c_{2}$ are equal.
5 Real-World Application
In this section, we discuss the application of the proposed model in developing a roadmap to address barriers to energy efficiency.
In the modern era, global energy consumption has increased at an alarming rate. This trend raises significant concerns regarding economic growth, environmental sustainability and the exhaustion of limited energy resources. A renewable energy source capable of fully replacing existing options has yet to be discovered. Therefore, the development and implementation of energy-efficient technologies have become imperative. However, such efforts are often hindered by implementation constraints. These barriers prevent societies from fully leveraging cost-effective and energy-friendly opportunities, making the study of obstacles to energy efficiency a critically important topic. Numerous studies have examined these barriers across various domains including industrial [9, 10], economic [25, 36], and environmental [23, 58] perspectives among others.
Buildings represent one of the highest energy-consuming sectors globally. In Europe, for instance, buildings account for a larger share of energy consumption than the industrial sector [63, 39]. Energy is consumed throughout a building’s entire life cycle including the production of raw materials, construction and demolition. The main reason behind this increasing energy demand and population growth and rising demand for building services [39]. The actual consumption level varies significantly from one building to another, depending mainly on local weather conditions, structure, and building characteristics [63]. Numerous studies have been conducted to predict energy consumption in buildings. Zhao et al. [63] comprehensively reviewed various engineering, statistical, and neural network models developed for this purpose. In a similar vein, Caldera et al. [11] analyzed statistical data to evaluate energy demand specifically for space heating in buildings.
This study aims to rank barriers to energy efficiency in buildings based on their criticality. Gupta et al. [21] identified $27$ such barriers and categorized them into six groups, as presented in Table 10. The pairwise comparisons between these categories and between barriers within the same category are adopted from their study.
The category weights, local and global weights of barriers, and the ranking of barriers based on these global weights are presented in Table 11. Our analysis, based on this model, shows that “scarcity of financial means” is the most critical barrier, while “contention” is the least critical barrier among all those ranked.
Table 10: Barriers to energy efficiency in buildings [21]
Category Driver
$c_{1}$: Economic barriers $c_{11}$: Scarcity of financial means
$c_{12}$: Absence of lucravity
$c_{13}$: Poor arbitrage
$c_{14}$: Inadequate monetary assessment
$c_{15}$: Limits in financial provisioning
$c_{2}$: Government barriers $c_{21}$: Lack of financial motivation
$c_{22}$: Bridles in hierarchical inspiration
    and functional harmony
$c_{23}$: Difference in plan of action
$c_{24}$: Inappropriate antecedencey
$c_{25}$: Lack of standards and references
$c_{26}$: Lack of strong authority
$c_{3}$: Knowledge and learning $c_{31}$: Lack of cognizance
    barriers $c_{32}$: Inexperient persons
$c_{33}$: Lack of information
$c_{4}$: Market related barriers $c_{41}$: Contention
$c_{42}$: Perceptual knowledge
$c_{43}$: Aspousal
$c_{44}$: Dubiety in demand
$c_{5}$: Organization and social $c_{51}$: Cognitive decision making approach
    barriers $c_{52}$: Lack of authority and jurisdiction
$c_{53}$: Ill defined vision
$c_{54}$: Torporpidity in process and practices
$c_{6}$: Technological barriers $c_{61}$: Incompatible technology
$c_{62}$: Process related risks
$c_{63}$: Lack of energy efficient materials
$c_{64}$: No feasibility study
$c_{65}$: Slow embodiment of new technology
Table 11: Category weights, barrier weights, and barrier ranking
Category Weight Driver Local weight Global weight Rank
$c_{1}$ $0.4107$ $c_{11}$ $0.4834$ $0.1985$ 11
$c_{12}$ $0.1196$ $0.0491$ 66
$c_{13}$ $0.0585$ $0.0240$ $13$
$c_{14}$ $0.1019$ $0.0419$ 99
$c_{15}$ $0.2366$ $0.0972$ 22
$\epsilon^{*}$ $0.2583$ - -
CR $0.0578$ - -
$c_{2}$ $0.2225$ $c_{21}$ $0.1440$ $0.0320$ $10$
$c_{22}$ $0.2025$ $0.0451$ 88
$c_{23}$ $0.4138$ $0.0920$ 33
$c_{24}$ $0.0501$ $0.0111$ 2020
$c_{25}$ $0.0873$ $0.0194$ $17$
$c_{26}$ $0.1023$ $0.0228$ $14$
$\epsilon^{*}$ $0.2583$ - -
CR $0.0578$ - -
$c_{3}$ $0.0989$ $c_{31}$ $0.7146$ $0.0707$ 44
$c_{32}$ $0.1$ $0.0099$ 2121
$c_{33}$ $0.1854$ $0.0183$ $18$
$\epsilon^{*}$ $0.1459$ - -
CR $0.0392$ - -
$c_{4}$ $0.0459$ $c_{41}$ $0.0713$ $0.0033$ $27$
$c_{42}$ $0.5809$ $0.0267$ $12$
$c_{43}$ $0.1443$ $0.0066$ $25$
$c_{44}$ $0.2035$ $0.0093$ $23$
$\epsilon^{*}$ $0.1459$ - -
CR $0.0326$ - -
$c_{5}$ $0.0847$ $c_{51}$ $0.2605$ $0.0221$ $15$
$c_{52}$ $0.1158$ $0.0098$ $22$
$c_{53}$ $0.5610$ $0.0475$ 77
$c_{54}$ $0.0627$ $0.0053$ $26$
$\epsilon^{*}$ $0.1538$ - -
CR $0.0294$ - -
$c_{6}$ $0.1373$ $c_{61}$ $0.4724$ $0.0649$ 55
$c_{62}$ $0.1580$ $0.0217$ $16$
$c_{63}$ $0.0975$ $0.0134$ 1919
$c_{64}$ $0.0528$ $0.0073$ $24$
$c_{65}$ $0.2193$ $0.0301$ $11$
$\epsilon^{*}$ $0.1538$ - -
CR $0.0294$ - -
$\epsilon^{*}$ $0.1538$ - - - -
CR $0.0294$ - - - -
@@ 6 Conclusions and Future Directions
The BWM has emerged as a powerful MCDM technique and is widely adopted in real-world applications. This study introduces an analytical framework for its nonlinear model, which is compatible with any scale and any number of DMs. The framework derives closed-form solutions for optimal interval-weights, CI, and CR. It also identifies the best optimal weight set from the collection of all optimal weight sets. Furthermore, it ranks the barriers to energy efficiency in buildings by their criticality. This research enhances the model’s efficiency in several key aspects. By eliminating the need for specialized optimization software, the framework significantly improves computational efficiency. It also rectifies existing inaccuracies in the calculation of CI. Additionally, it transforms CR into an input-based consistency indicator, capable of providing immediate feedback to DMs. The framework also modifies the original optimization model for cases with multiple best or worst criteria to select the optimal weight sets which assign equal weights to all best criteria and equal weights to all worst criteria.
A promising future research direction emerging from this work is to derive analytical expressions for optimal weights, CI, and CR for models such as the Euclidean BWM [28] and the $\alpha$-FBWM [40] for which such closed-form solutions have not yet been established. Developing such frameworks would significantly enhance the applicability and efficiency of these methods.